# The Sparse Tsetlin Machine: Sparse Representation with Active Literals

### v1

@@ Abstract
Tsetlin Machine, Natural Language Processing, Interpretable AI , Sparse, Sparse Data Processing, Sparse Clauses
@@ I Introduction
If condition A and not condition B then Result 1.
The TM yields competitive results compared to other classification methods when performing benchmarks, and the hardware-near approach reduces the energy footprint and minimizes memory impact during training and inference [2]. TMs encompass a diverse range of machine learning paradigms, including classification[1], convolution[3], regression[4], and reinforcement learning [5]. Recent research has harnessed TMs for various applications, such as sentiment analysis [6], fake news detection [7], and novelty detection [8][9]. TMs have also been applicable in the healthcare sector [10].
@@
This paper presents the Sparse Tsetlin Machine (STM), a novel TM that takes advantage of sparsity by introducing Active Literals (AL). The AL is a gatekeeper at a class level that introduces literals and sub-patterns from the data into clauses via frequent pattern mining. This enables effective sparse data processing. The STM will only distribute resources to discriminative literals, ensuring that only literals that actively influence predictions are considered. Novel to the STM is the ability to initialize the TM memory with empty clauses, with no record of TAs at the beginning of training. Then, through feedback, the AL introduces literals and sub-patters contained in the data into memory. Further, a lower-bound threshold creates a minimum TA state allowed in clauses. This establishes an allowed state spectrum where TAs can fluctuate. If a state is reduced below the allowed lower bound, the TA is removed from the clause immediately, making the memory smaller and patterns contained in clauses robust and concise. The product results in a new efficient sparse memory space for the STM that efficiently performs memory feedback, distributing resources solely to literals that actively contribute to the current data representation, dramatically increasing the processing capabilities of sparse data. The STM is extensively evaluated on a vast amount of well-known datasets. When processed for TM training, these datasets become vastly sparse and are ideal for confirming the STM’s ability to process sparse data while achieving competitive performance efficiently. Furthermore, the compact memory usage of the STM allows traditional TM training on large-scale datasets, such as big text corpora, a previously unattainable capability.
@@ III Sparse Tsetlin Machine
A fundamental new input space is required to process sparse data efficiently. The STM operates under the assumption of the sparse data structure, Compressed Sparse Row (CSR). This structure depicts only the active, non-zero values in the data, compressing the input space. However, the STM goes one step further by discarding all negated features. Traditional TM training assumes the expansion of literals to $2o$ features per instance, with literals $x_{k}$ for $o+1\leq k\leq 2o$ exclusively representing negated features, effectively doubling the data size at every instance. On the other hand, lossless compression can be leveraged by discarding negated features and employing CSR. Assuming the data dimensions are known beforehand, the compression is lossless. Since if a literal is active in the input, it is implicitly understood that its negated form is not present. Conversely, if a literal is not active in the input, it is understood to be in its negated form. This understanding enables discarding all negated representations, avoiding expansion to $2o$ features, and making the input space more compact, simpler, and efficient. Verifying the presence or absence of a literal can be concluded if corresponding negated features are in effect. Therefore, the STM input space is more diminutive and compact by leveraging sparsity using an accurate sparse representation by discarding zero values.
@@ III-B Sparse Memory Space
The sparse clause structure is defined with lists. Novel to the STM, with the introduction of the AL, is the ability for sparse clauses to start empty without any record of existing literals or states. Then, through feedback, literals are introduced to clauses from frequently observed patterns in the AL, populating the clauses with only the essential literals observed from the data. There are two main advantages to this approach. Firstly, it has a smaller memory impact because literals are introduced gradually instead of initializing a subset or the entire memory. This helps avoid memory limitations by allocating space only for features relevant to inference. Secondly, this approach makes memory more flexible, as literals can be introduced without requiring pre-defined representations. This allows the introduction of new features on the go.
@@
TAs not present in the clauses, either discarded or not introduced, are not defined, and there is no record of their state or action. With its dynamic introduction and discard of TAs during training, the new structure enables more efficient training by only distributing resources and feedback to literals found to be active by the AL and introduced to clauses.
@@ III-C Updating the memory space
The learning scheme of the STM follows that of the CoTM closely, described in Section II-A. In addition, Focused Negative Sampling (FNS) is utilized to better distinguish between classes by forcing the closest classes further away from each other [12]. Feedback only differs in updating the AL, which occurs exclusively in feedback types Ia and II.
@@
During the training phase, the STM memory space gets populated with literals from the AL. For literals introduced into memory, feedback is distributed traditionally as the CoTM scheme depicts. A strict limit to the maximum number of literals allowed in a clause is defined to prevent clauses from being overpopulated. Defined as $p$, the maximum clause size determines the number of literals introduced to any clause at any given time. Suppose a clause is at its maximum capacity; a literal must be decarded by being reduced below the allowed threshold before allowing new introductions.
@@ IV Empirical results
In this chapter, we present the empirical results of the STM. We conducted experiments to confirm the STM’s ability to effectively leverage sparse data for accurate predictions. We have organized the results into three sections, each focusing on a different aspect of performance. Firstly, we present a series of experiments conducted on various datasets, each with varying degrees of sparsity. We recorded the performance for each case. Secondly, we examine the ability to maintain efficiency across an increasingly sparse problem. We discuss how the STM’s performance evolves as the size of the dataset increases and highlight the ability to maintain high performance and efficiency when dealing with an increasingly sparse problem. Lastly, we employed the STM in a large-scale text corpus application, demonstrating the STM’s ability to work in compressed input space and, thus, significantly lower memory usage.
@@ IV-A Experiments details
[…]
@@ IV-C Results
An experiment was conducted using the SUBJ dataset to evaluate the STM’s proficiency in handling sparse data. The experiment involved increasing the vocabulary size, denoted as $\hat{V}$ in $(n\times\hat{V})$, to monitor the performance on increasingly sparse problems. As the vocabulary size increases, the data becomes more complex and high-dimensional, challenging the conventional characteristics of TM training that become inefficient with larger input spaces. The experiment began with a vocabulary size of $\hat{V}=2500$ and was gradually increased by 500 until it reached $\hat{V}=10000$. The results of the experiment are presented in Fig. 5 and 6.
@@ IV-D Deploying STM on large scale text corpus
The STM’s excellent processing of sparse data opens up new possibilities for large-scale text corpus applications. Its efficient processing of sparse data and diminutive memory impact make it an ideal solution for handling large volumes of textual data. The Amazon Review All data bank is a large-scale text corpus accumulating $233.1$ million in individual textual documents, representing product reviews from 5 classification labels. We conducted five experiments to explore the STM’s scalability by randomly sampling a subset of $n$ documents from the text corpus. Initially, we start with a small subset, gradually increasing the number of documents $n$ and vocabulary sizes $\hat{V}$ to stress the STM’s ability. The experiment was conducted using traditional TM training without batching.

### v2

@@ Abstract
Tsetlin Machine, Natural Language Processing, Interpretable AI , Sparse Data Processing, Sparse Clause Representation
@@ I Introduction
If Condition A and not Condition B then Result C.
The TM yields competitive results compared to other classification methods when performing benchmarks, and the hardware-near approach reduces energy footprint and minimizes memory impact during training and inference [2]. TMs encompass a diverse range of machine learning paradigms, including classification[1], convolution[3], regression[4], and reinforcement learning [5]. Recent research has harnessed TMs for various applications, such as sentiment analysis [6], fake news detection [7], and novelty detection [8][9]. TMs have also been applicable in the healthcare sector [10].
@@
This paper presents the Sparse Tsetlin Machine (STM), a novel TM that introduces Active Literals (AL) to leverage sparsity. The AL acts as a gatekeeper at a class level, introducing literals and sub-patterns from the data into clauses via frequent pattern mining. This enables effective sparse data processing. The STM will only distribute resources to discriminative literals, ensuring that only literals that actively influence predictions are considered. Novel to the STM is the ability to initialize the TM memory with empty clauses that contain no record of TAs at the beginning of training. Then, through feedback, the AL introduces literals and sub-patters contained in the data into memory. Further, a lower-bound threshold creates a minimum TA state allowed in clauses. This establishes an allowed state spectrum where TAs can fluctuate. If a state is reduced below the allowed lower bound, the TA is removed from the clause immediately, making the memory smaller and patterns contained in clauses robust and concise. This coalesces in a new efficient sparse memory space for the STM that efficiently performs memory feedback, distributing resources solely to literals that actively contribute to the current data representation, dramatically increasing the processing capabilities of sparse data. The STM is extensively evaluated on a vast amount of well-known datasets. When processed for TM training, these datasets become vastly sparse and are ideal for confirming the STM’s ability to process sparse data while achieving competitive performance efficiently. Furthermore, the compact memory usage of the STM permits traditional TM training on large-scale datasets, such as big text corpora, a previously unattainable capability.
@@ III Sparse Tsetlin Machine
A fundamental new input space is required to process sparse data efficiently. The STM operates under the assumption of the sparse data structure, Compressed Sparse Row (CSR). This structure depicts only the active, non-zero values in the data, compressing the input space. However, the STM goes one step further by discarding all negated features. Traditional TM training assumes the expansion of literals to $2o$ features per instance, with literals $x_{k}$ for $o+1\leq k\leq 2o$ exclusively representing negated features, effectively doubling the data size at every instance. On the other hand, assuming the data dimensions are known beforehand, lossless compression can be leveraged by discarding negated features and employing CSR. In cases where literals are present in the input, their negated form is implicitly assumed to be absent. Similarly, if a literal is not present in the input, it is understood to be in its negated form. This understanding enables discarding all negated representations, avoiding expansion to $2o$ features, and making the input space more compact, simpler, and efficient. The STM verifies the presence or absence of literals and concludes whether corresponding negated features are in effect. Therefore, input space is more diminutive and compact by leveraging sparsity using an accurate sparse representation by discarding zero values.
@@ III-B Sparse Memory Space
The sparse clause structure is defined with lists. Novel to the STM, with the introduction of the AL, is the ability for sparse clauses to start empty without any record of existing literals or states. Then, through feedback, literals are introduced to clauses from frequently observed patterns in the AL, populating the clauses with only the essential literals observed from the data. There are two main advantages to this approach. Firstly, instead of initializing a subset [2] or the entire memory all at once [1], literals are introduced gradually, allocating space only for the features relevant to inference. This not only helps to avoid memory limitations but also ensures that resources are used efficiently. Secondly, this approach makes memory more flexible, as literals can be introduced without requiring pre-defined representations, allowing for the introduction of new features on the go.
@@
TAs not present in the clauses, either discarded or not introduced, are not defined, and there is no record of their state or action. With its dynamic introduction and discarding of TAs during training, the new structure enables more efficient training by only distributing feedback to literals deliberately introduced into memory.
@@ III-C Updating the memory space
The learning scheme of the STM follows that of the CoTM closely, described in Section II-A. In addition, Focused Negative Sampling (FNS) is utilized to better distinguish between classes by forcing similar classes further away from each other [12]. Feedback only differs in updating the AL, which occurs exclusively in feedback types Ia and II.
@@
During the training phase, the STM memory space gets populated with literals from the AL. For literals introduced into memory, feedback is distributed traditionally as the CoTM scheme depicts. To prevent a clause from getting overpopulated, a strict limit for the maximum number of literals allowed in clauses is enforced. Defined as $p$, the maximum clause size determines the number of literals introduced to any clause at any given time. Suppose a clause is at its maximum capacity; a literal must be discarded by being reduced below the allowed threshold before allowing new introductions.
@@ IV Empirical results
In this chapter, we present the empirical results of the STM. We conducted experiments to confirm the STM’s ability to effectively leverage sparse data for accurate predictions. We have organized the results into three sections, each focusing on a different performance aspect. Firstly, we present a series of experiments conducted on various datasets, each with varying degrees of sparsity. We recorded the performance for each case. Secondly, we examine the ability to maintain efficiency across an increasingly sparse problem. We discuss how the STM’s performance evolves as the size of the dataset increases and highlight the ability to maintain high performance and efficiency when presented with increasingly sparse problems. Lastly, we employed the STM in a large-scale text corpus application, demonstrating the STM’s ability to work in compressed input space and, thus, significantly lower memory usage.
@@ IV-A Experiments details
We initialized the STM with the following parameters: Number of clauses ($N$), voting margin ($T$), specificity ($s$), active literal size ($a$), lower-state threshold ($t$), and max clause size ($p$). We measured sparsity as the percentage of non-zero values in any data representation. For the first experiment, we tuned the STM for each dataset. In the second experiment, we increased the vocabulary size of an NLP problem. This was done to expand the data while producing increasingly sparse representations. Finally, we performed five experiments on a large-text corpus. We iteratively increased the total number of documents and vocabulary sizes. Unless otherwise noted, all experiments were conducted on an Intel i5-13600KF @ $3.5GHz$, running single-threaded on Ubuntu 20.04 LTS.
@@ IV-C Results
An experiment was conducted using the SUBJ dataset to verify the STM’s sparse data processing capabilities. The experiment involved increasing the vocabulary size, denoted as $\hat{V}$ in $(n\times\hat{V})$, to monitor the performance on increasingly sparse problems. As the vocabulary size increases, the data becomes more complex and high-dimensional, challenging the conventional characteristic of TM training becoming inefficient with large input spaces. The experiment began with a vocabulary size of $\hat{V}=2500$ and was gradually increased by 500 until it reached $\hat{V}=10000$. The results of the experiment are presented in Fig. 5 and 6.
@@ IV-D Deploying STM on large scale text corpus
The STM’s excellent processing of sparse data opens up new possibilities for large-scale text corpus applications. Its efficient processing of sparse data and diminutive memory impact make it an ideal solution for handling large volumes of textual data. The Amazon Review All data bank is a large-scale text corpus accumulating $233.1$ million in individual textual documents, representing product reviews with 5 classification labels. We conducted five experiments to explore the STM’s scalability by randomly sampling a subset of $n$ documents from the text corpus. Initially, we start with a small subset, gradually increasing the number of documents $n$ and vocabulary sizes $\hat{V}$ to stress the STM. The experiment was conducted using traditional TM training without batching.